Circular No. 6069

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

JUPITER AND COMET 1993e M. K. Bird and O. Funke, University of Bonn; J. Neidhoefer, Max-

Planck-Institut für Radioastronomie (MPIR), Bonn; and I. de Pater, Department of Astronomy, University of California at Berkeley, communicate: “Radio observations of Jupiter were performed at the Effelsberg Telescope before, during, and after the July 16–22 impacts of comet 1993e. Measurements of the four Stokes parameters, yielding full-polarization characteristics of the received signal, were recorded at three wavelengths: 2.8, 6, and 11 cm. Generally consistent with similar radio observations reported by Klein and Gulkis (IAUC 6042), Galopeau et al. (IAUC 6045), and Smits et al. (IAUC 6047), Jupiter’s radio brightness was found to increase monotonically during the impact week at all three wavelengths. The increase from July 16/17 to 21/22 was computed at a representative point in the beaming curve when the system-III central-meridian longitude was 290^{o} (near the absolute peak). The absolute (relative) increases in total flux, referred to the standard distance of 4.04 AU, were 1.2 ± 0.1 Jy (18 percent) at 11 cm and 1.2 ± 0.2 Jy (13 percent) at 6 cm. A preliminary analysis of the 11-cm observations shows an increase during the impact week of ∼ 2 percent in the relative linear polarization, but no immediately-apparent changes in the polarization (position) angle or circular polarization. Post-impact observations taken on Aug. 12/13 indicate a slight decrease (∼ 0.2 Jy) in total flux with respect to the level recorded on July 21/22. Jupiter’s radio brightness observed at the most recent Effelsberg observations on Aug. 26/27 remained unchanged from the value on Aug. 12/13. An impact-induced relative increase in flux at 2.8 cm is apparent, but further data processing is required before a quantitative statement can be made.”

POSSIBLE GRAVITATIONAL MICROLENSING EVENT S. Benetti, L. Pasquini, and R. West, European Southern Observatory

(ESO); and V. de Lapparent, Institut d’Astrophysique de Paris, report that a provisional inspection of a wavelength-calibrated spectrum (391–880 nm) obtained on Sept. 2.07 UT with the ESO 3.5-m New Technology Telescope (+ EMMI) of the possible MACHO (cf. IAUC 6068) reveals no emission lines, a comparatively smooth continuum, and the following main absorption features (relative strengths indicated): G-band (medium); Hβ (weak); Mg I b triplet (strong); Fe I/Cr I complex (578.0-579.0 nm; medium); Na I D doublet (very strong); Hα (medium). The spectrum is consistent with that of a normal K0–K2 III–V star.

1994 September 2 Daniel W. E. Green